# Dual topological insulator with mirror-symmetry-protected helical edge states

### v1

@@ Abstract
Dual topological insulators (DTIs) are simultaneously protected by time-reversal and crystal symmetries, representing advantageous alternatives to conventional topological insulators. By combining ab initio calculations and the $\bm{k}\cdot\bm{p}$ approach, here we investigate the electronic band structure of a Na2CdSn tri-atomic layer and derive a low-energy $4\times 4$ effective model consistent with all the symmetries of this material class. We obtain the effective Hamiltonian using the Löwdin perturbation theory, the folding down technique, and the theory of invariants, and determine its parameters by fitting our analytical dispersion relations to those of ab initio calculations. We then calculate the bulk topological invariants of the system and show that the Na2CdSn tri-atomic layer is a giant-gap (hundreds of meV) quasi-2D DTI characterized by both spin and mirror Chern numbers $-2$. In agreement with the bulk-boundary correspondence theorem, we find that a finite-width strip of Na2CdSn possesses two pairs of counter-propagating helical edge states per interface. We obtain analytical expressions for the edge states energy dispersions and wave functions, which are shown to agree with our numerical calculations. Our work opens a new avenue for further studies of Na2CdSn as a potential DTI candidate with room-temperature applications in areas of technological interest, such as nanoelectronics and spintronics.
@@ I Introduction
Since the theoretical proposal and experimental observation of 2D [1, 2, 3, 4, 5, 6] and 3D [7, 8, 9, 10, 11, 12] topological insulators (TIs), these materials have been the subject of substantial research in condensed matter physics and related areas of science [11, 13, 14, 15, 16, 17]. Further investigations soon resulted in the discovery of another important class of topological materials named topological crystalline insulators (TCIs) [18, 19, 20, 21, 22]. Both TIs and TCIs host mid-gap boundary states related to a bulk topological invariant via the bulk-boundary correspondence [23, 24]. The TI phase can be characterized by a non-trivial spin Chern number (in 2D) [25, 26, 27, 28] or $Z_{2}$ invariant [1, 3, 8], whereas the characterization of a TCI phase depends on the particular material and involves at least one crystal symmetry, e.g., inversion, rotation, or the mirror operation (it may or may not involve TRS). Well-known examples are TCIs characterized by the mirror Chern number [29, 30, 20, 31] and the $Z_{2}$ invariant defined in terms of the product of time reversal with a $C_{4}$ rotation [18].
@@
The interplay between time-reversal and crystalline symmetries on the topology of 3D materials was first investigated in Bi1-xSbx  [32, 33] and Bi2Te3 [34] alloys, which were predicted to be not only $Z_{2}$ TIs, but also TCIs characterized by a non-trivial mirror Chern number. Such a dual topological character defines Bi1-xSbx, Bi2Te3 and other materials [35, 36, 37, 38, 39] as dual topological insulators (DTIs), considered more robust against external perturbations than each constituent phase [40, 36, 34]. For example, it has been shown that to open a gap in the Dirac cone describing the $(111)$ surface states of Bi2Te3 one must break time reversal and mirror symmetries simultaneously [34].
@@
Dual topological insulators in 2D have been much less studied than their 3D counterparts. The literature is nearly limited to theoretical studies of the Na3Bi single layer system [41, 42, 43]. More recently, the rectangular Bi bilayer [40] and the hexagonal IrO [44] have also been proposed. Up to now, to the best of our knowledge, these have not yet been experimentally realized.
In this work, we explore an overlooked DTI constituted by a tri-atomic layer of the ternary sodium compounds Na2XY (X=Mg,Cd; Y=Pb,Sn) which can, potentially, be obtained by mechanical exfoliation of 3D van der Waals Dirac semimetals [45] already synthesized experimentally [46, 47]. We derive a $\bm{k}\cdot\bm{p}$ Hamiltonian for this class of materials (provided that the ordering of the energy bands around the Fermi level is the same as that for Na2CdSn) and analyze the bulk topology and the corresponding mirror-symmetry-protected helical edge states of its Na2CdSn representative. By carrying out density functional theory (DFT) calculations, we first obtain the electronic band structure of Na2CdSn, including the irreducible representations (IRREPs), orbital composition and spin texture of the energy bands (Fig. 1). Then, we derive our $4\times 4$ model Hamiltonian for the Na2XY material class by employing group theory analysis and three techniques well-established in the $\bm{k}\cdot\bm{p}$ literature, namely, the Löwdin perturbation theory [48, 49], the folding down approach [50, 51, 52], and the theory of invariants [53, 49].
To obtain the parameters of our model, we diagonalize our effective Hamiltonian and fit its analytical energy dispersions to our DFT results. We then investigate the “orbital” (pseudospin) composition and spin texture of the effective band structure (Fig. 2). As a consequence of mirror symmetry [54] (see Appendix A), the energy bands are spin-polarized along the $z$ direction, in agreement with the DFT calculations (Fig. 1). Furthermore, by regularizing our model and numerically evaluating the Chern number [55] for each spin sector, we show that Na2CdSn is a DTI with the TI and TCI phases characterized by spin and mirror Chern numbers $-2$, respectively (Fig. 3).
We investigate the DTI spectrum for two finite systems: a semi-infinite plane (vacuum/DTI) and a ribbon geometry (vacuum/DTI/vacuum). For the semi-infinite plane, we employ the envelope function approach to analytically derive two pairs of counter-propagating topological edge states related by TRS (Kramers pairs) whose dispersion relations depend quadratically on the crystal momentum (Fig. 4). For the ribbon geometry, we use the finite difference method to numerically obtain its electronic band structure and identify the existence of four pairs of counter-propagating edge states, two pairs located on each interface of the system. The states localized at opposite sides of the ribbon are quasi-degenerate, with a small spin-splitting caused by the bulk inversion asymmetry (BIA). Our results for both geometries are in agreement with the bulk-boundary correspondence theorem [23, 24].
The extra robustness against external perturbations [40, 36, 34] positions DTIs as promising materials for applications in nanotechnology. In particular, the Na2CdSn system investigated here is even more compelling. According to our DFT calculations, its expected bulk energy gap, $E_{g}\approx 234.8\,$meV, is about one order of magnitude larger than those of 2D TIs based on HgTe/CdTe, InAs/GaSb or InAs0.85Bi0.15/AlSb quantum wells, for which $E_{g}$ $\lesssim 30\,$meV [6, 56, 57, 58]. Such features showcase this ternary compound as an outstanding candidate for the study of room-temperature topological effects and the development of new nanoelectronic [59, 60, 61], spintronic [62, 63, 64], thermoelectric [65, 66, 67] and optical [68, 69] devices.
We note that a study of the Na2XY tri-atomic layers was performed in Ref. [70]. The authors propose an effective model Hamiltonian  [^1] The authors of Ref. [70] provide very limited information about the basis set used in their work. This precluded us from verifying whether their reported effective Hamiltonian respects all the symmetries of the system – in particular time-reversal symmetry. that can be immediately seen to differ from ours by i) the lack of terms proportional to $(k_{x}\pm ik_{y})^{2}$ in the off-diagonal matrix elements and ii) the presence of terms proportional to $k_{x}\pm ik_{y}$ in the secondary diagonal, where $\bm{k}=(k_{x},k_{y})$ represents the crystal momentum. It is important to emphasize that we consistently kept terms up to second order in $\bm{k}$ in the perturbation theory.
As we discuss in Sec. IV, the terms proportional to $(k_{x}\pm ik_{y})^{2}$ play a crucial role in the calculation of the spin and mirror Chern numbers, which are both equal to $-2$. Moreover, they give rise to the quadratic energy dispersions of the corresponding topological edge states. The topological invariants obtained from our analytical model differ significantly from the spin and mirror Chern numbers $-1$ reported in Ref. [70]. Following the bulk-boundary correspondence theorem, we predict the existence of two pairs of helical edge states at each boundary of a finite Na2XY system, whereas Ref. [70] reports only a single pair.
To ensure the accuracy of our model, we first performed the derivation of the $\bm{k}\cdot\bm{p}$ Hamiltonian using the Löwdin perturbation theory, and separately, using the folding down approach. We also derived our effective Hamiltonian using the theory of invariants. The results obtained are in complete agreement with each other, and differ from the model proposed in Ref. [70]. In addition, our investigation here goes beyond that of Ref. [70] by providing effective $\bm{k}\cdot\bm{p}$ parameters for the Na2CdSn DTI and deriving analytical expressions for both the energy dispersion and wave functions of the edge states arising in this material.
The article is organized as follows. In Sec. II, we present the crystal structure and DFT calculations for Na2CdSn tri-atomic layers. Sec. III is devoted to the $\bm{k}\cdot\bm{p}$ model obtained via the Löwdin perturbation theory. We also present our results regarding the fitting of the effective model to the energy bands obtained via DFT. In Sec. IV we analyze the topology of Na2CdSn and, in Sec. V, we calculate the analytical dispersion relations for the topological edge states. We also present our numerical study of the topological edge states in a ribbon geometry. We summarize our findings in Sec. VI.
II Na2CdSn tri-atomic layer: Crystal structure and DFT results
The Na2CdSn tri-atomic layer is constituted by Cd and Sn atoms forming a co-planar hexagonal lattice, while the sodium (Na) atoms lie above and below the center of each hexagon, forming a quasi-2D honeycomb material (see Fig. 1). This quasi-2D structure can be found as the constituting stacking block of the 3D Na2CdSn Dirac semimetal, already synthesized experimentally [46, 47].
@@ II Na2CdSn tri-atomic layer: Crystal structure and DFT results
In Fig. 1-a), we show a top view of the Na2CdSn. The coordinate system is chosen such that the $z$-axis points out of the plane of the figure. The primitive lattice vectors read $\bm{a}_{1}=\frac{a}{2}\left(\sqrt{3},1\right)$ and $\bm{a}_{2}=\left(0,a\right)$, where $a$ is the lattice parameter. In Fig. 1-b), we show a side view of the crystal structure with the observer looking from the positive $x$-axis. From this angle, the $x$-axis points out of the plane of the figure, and we can see more clearly the position of the Na atoms relative to Cd and Sn. The reciprocal lattice vectors can be shown to be $\bm{b}_{1}=\left(\frac{4\pi}{\sqrt{3}a},0\right)$ and $\bm{b}_{2}=\frac{2\pi}{\sqrt{3}a}\left(-1,\sqrt{3}\right)$. The first Brillouin zone (BZ) in this coordinate system is shown in Fig. 1-c), along with the high-symmetry points $\bm{\Gamma}=\left(0,0\right)$, $\bm{M}=\left(\frac{2\pi}{\sqrt{3}a},0\right)$ and $\bm{K}=\frac{2\pi}{3a}\left(\sqrt{3},1\right)$ [72].
@@
The crystal structure of the 3D bulk compound belongs to the non-symmorphic space group $D_{6h}^{4}$ ($P6_{3}/mmc$), with point group $D_{6h}$ ($6/mmm$) [73, 47, 45, 74, 75]. Our DFT results show that the Na2CdSn belong to the $D_{3h}^{1}$ ($P\bar{6}m2$) space group, whose crystallographic point group is the $D_{3h}$ ($\bar{6}m2$) [73, 47]. The corresponding set of generators are the $C_{3}$ rotation about the $z$-axis ($C_{3z}$), $C_{2}$ rotation about the $x$-axis ($C_{2x}$) and the mirror plane reflection about the $xy$-plane ($\sigma_{h}$) [see character table in the Supplemental Material (SM) [76]].
@@
We investigate the electronic structure of the Na2CdSn using DFT as implemented in the WIEN2k package [77], a full potential all-electron code employing the linearized augmented plane wave plus local orbitals method. We use the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [78], a Monkhorst-Pack k-grid of 15$\times$15 and self-consistent convergence criteria of 10-6 $e$ (elementary charge) for the charge and 10-6 Ry for the energy. We consider a core–valence separation energy of $-6$ Ry, atomic spheres with orbital quantum numbers up to 10 and the plane-wave cutoff multiplied by the smallest atomic radii is set to 8. In the absence of spin-orbit coupling (SOC), we optimize the structural parameters and obtain an in-plane lattice parameter $a=4.978\,$Å and a thickness (vertical distance between Na atoms) of 3.3323.3323.332\,3.332Å (consistent with the lattice parameters of the 3D bulk compound [45, 46]). In the presence of SOC, core electrons are considered fully relativistically whereas valence electrons are treated in a second variational step [79], with the scalar-relativistic wave functions calculated in an energy window of $-10$ to $4$ Ry. We consider a vacuum region of 202020\,20Å.
@@
In Figs. 1-d) and 1-e), we show the electronic band structure and the orbital composition, respectively, obtained via DFT calculations for Na2CdSn in the absence of SOC. The Fermi level ($E_{F}$) lies at zero energy and the blue dashed circle indicates the low-energy region we are interested in. The band structure is that of a semimetal with two $\Gamma_{6}$ bands fourfold degenerate at the $\Gamma$ point (accounting for the twofold spin-degeneracy of each band). The closest remote bands belong to the $\Gamma_{1}$ and $\Gamma_{4}$ irreducible representations (IRREPs) [76]. We note that the $\Gamma_{6}$ bands are dominated by the $p_{x}$ and $p_{y}$ orbitals of the Cd and Sn atoms [dark blue in 1-e)], while the $\Gamma_{1}$ bands are dominated by $s$ orbitals (yellow) with a minor contribution from $p_{x}$ and $p_{y}$ (for more details see SM [76]).
@@
In the presence of SOC, Fig. 1-f), the degeneracy at the $\Gamma$ point is lifted and an energy gap, $E_{g}=234.8\,$meV, opens up. The material then becomes a semiconductor with topmost valence bands belonging to the $\Gamma_{7}$ and $\Gamma_{8}$ IRREPs, while the lowest conduction bands belong to $\Gamma_{7}$ and $\Gamma_{9}$. The next remote conduction band belongs to the $\Gamma_{8}$ IRREP. All IRREPs are obtained from the total wave function.
@@
In Fig. 1-g), we show the expectation value of the $z$-component of spin (color-coded), $\langle\hat{S}_{z}\rangle$ [$\langle\hat{S}_{z}\rangle>0$ ($\langle\hat{S}_{z}\rangle<0$) red (blue)]. We note that along the $\Gamma-M$ high-symmetry line ($\Sigma$), all energy bands are doubly degenerate. This high-symmetry line belongs to the point group $C_{2v}$, whose double group has only one IRREP – the two-dimensional $\Sigma_{5}$ [80]. The spin-up and -down components of the electron wave functions transform according to the two basis functions of $\Sigma_{5}$. Therefore, all energy bands must be spin-degenerate along this direction. Since any linear combination of the spin-degenerate states is also an eigenstate of the Hamiltonian, $\langle\hat{S}_{z}\rangle$ is not univocally defined along $\Gamma-M$. Hence we represent the degenerate bands using the black color.
@@
The $\Gamma-K$ ($T$) and $M-K$ ($T^{\prime}$) lines, on the other hand, belong to the point group $C_{s}$, which consists of only two elements: the identity and the mirror symmetry $\sigma_{h}$. Its double group representation has two one-dimensional IRREPs, $T_{3}$ ($T^{\prime}_{3}$) and $T_{4}$ ($T^{\prime}_{4}$) for spin-up and -down states, respectively, along the $T$ ($T^{\prime}$) line. Thus the energy bands are spin-split along these directions. The spin expectation values of all eigenstates out of the $\Gamma-M$ line are given by $\langle\hat{S}_{z}\rangle\approx\pm\hbar/2$ and $\langle\hat{S}_{x}\rangle=\langle\hat{S}_{y}\rangle=0$, as required by mirror symmetry about the $z$ direction (see Appendix A) [54].

### v2

# Dual topological insulator with mirror symmetry protected helical edge states

@@ Abstract
Dual topological insulators (DTIs) are simultaneously protected by time-reversal and crystal symmetries, representing advantageous alternatives to conventional topological insulators. By combining ab initio calculations and the $\bm{k}\cdot\bm{p}$ approach, here, we investigate the electronic band structure of a Na2CdSn triatomic layer and derive a low-energy $4\times 4$ effective model consistent with all the symmetries of this material class. We obtain the effective Hamiltonian using the Löwdin perturbation theory, the folding-down technique, and the theory of invariants and determine its parameters by fitting our analytical dispersion relations to those of ab initio calculations. We then calculate the bulk topological invariants of the system and show that the Na2CdSn triatomic layer is a giant-gap (hundreds of millielectronvolts) quasi-two-dimensional DTI characterized by both spin and mirror Chern numbers $-2$. In agreement with the bulk-boundary correspondence theorem, we find that a finite-width strip of Na2CdSn possesses two pairs of counterpropagating helical edge states per interface. We obtain analytical expressions for the edge state energy dispersions and wave functions, which are shown to agree with our numerical calculations. Our work opens an avenue for further studies of Na2CdSn as a potential DTI candidate with room-temperature applications in areas of technological interest, such as nanoelectronics and spintronics.
@@ I Introduction
Since the theoretical proposal and experimental observation of two-dimensional (2D) [1, 2, 3, 4, 5, 6] and three-dimensional (3D) [7, 8, 9, 10, 11, 12] topological insulators (TIs), these materials have been the subject of substantial research in condensed matter physics and related areas of science [11, 13, 14, 15, 16, 17]. Further investigations soon resulted in the discovery of another important class of topological materials named topological crystalline insulators (TCIs) [18, 19, 20, 21, 22]. Both TIs and TCIs host midgap boundary states related to a bulk topological invariant via the bulk-boundary correspondence [23, 24]. The TI phase can be characterized by a nontrivial spin Chern number (in 2D) [25, 26, 27, 28] or $Z_{2}$ invariant [1, 3, 8], whereas the characterization of a TCI phase depends on the material and involves at least one crystal symmetry, e.g., inversion, rotation, or the mirror operation [it may or may not involve time-reversal symmetry (TRS)]. Well-known examples are TCIs characterized by the mirror Chern number [29, 30, 20, 31] and the $Z_{2}$ invariant defined in terms of the product of time reversal with a $C_{4}$ rotation [18].
@@
The interplay between time-reversal and crystalline symmetries on the topology of 3D materials was investigated in Bi1-xSbx  [32, 33] and Bi2Te3 [34] alloys, which were predicted to be not only $Z_{2}$ TIs but also TCIs characterized by a nontrivial mirror Chern number. Such a dual topological character defines Bi1-xSbx, Bi2Te3, and other materials [35, 36, 37, 38, 39] as dual TIs (DTIs), considered more robust against external perturbations than each constituent phase [40, 36, 34]. For example, it has been shown that, to open a gap in the Dirac cone describing the $(111)$ surface states of Bi2Te3, one must break time-reversal and mirror symmetries simultaneously [34].
@@
DTIs in 2D have been much less studied than their 3D counterparts. The literature is nearly limited to theoretical studies of the Na3Bi single-layer system [41, 42, 43]. More recently, the rectangular Bi bilayer [40] and the hexagonal IrO [44] have also been proposed. Up to now, to the best of our knowledge, these have not yet been experimentally realized.
@@
To obtain the parameters of our model, we diagonalize our effective Hamiltonian and fit its analytical energy dispersions to our DFT results. We then investigate the orbital (pseudospin) composition and spin texture of the effective band structure (Fig. 2). As a consequence of mirror symmetry [54] (see Appendix A), the energy bands are spin polarized along the $z$ direction, in agreement with the DFT calculations (Fig. 1). Furthermore, by regularizing our model and numerically evaluating the Chern number [55] for each spin sector, we show that Na2CdSn is a DTI with the TI and TCI phases characterized by spin and mirror Chern numbers $-2$, respectively (Fig. 3).
We investigate the DTI spectrum for two finite systems: a semi-infinite plane (vacuum/DTI) and a ribbon geometry (vacuum/DTI/vacuum). For the semi-infinite plane, we employ the envelope function approach to analytically derive two pairs of counterpropagating topological edge states related by TRS (Kramers pairs) whose dispersion relations depend quadratically on the crystal momentum (Fig. 4). For the ribbon geometry, we use the finite difference method to numerically obtain its electronic band structure and identify the existence of four pairs of counterpropagating edge states, two pairs located on each interface of the system (Fig. 5). The states localized at opposite sides of the ribbon are quasidegenerate, with a small spin splitting caused by the bulk inversion asymmetry (BIA). Our results for both geometries are in agreement with the bulk-boundary correspondence theorem [23, 24].
The extra robustness against external perturbations [40, 36, 34] positions DTIs as promising materials for applications in nanotechnology. The Na2CdSn system investigated here is even more compelling. According to our DFT calculations, its expected bulk energy gap, $E_{g}\approx 234.8\,$meV, is $\sim$1 order of magnitude larger than those of 2D TIs based on HgTe/CdTe, InAs/GaSb or InAs0.85Bi0.15/AlSb quantum wells, for which $E_{g}$ $\lesssim 30\,$meV [6, 56, 57, 58]. Such features showcase this ternary compound as an outstanding candidate for the study of room-temperature topological effects and the development of nanoelectronic [59, 60, 61], spintronic [62, 63, 64], thermoelectric [65, 66, 67], and optical [68, 69] devices.
[…]
To ensure the accuracy of our model, we first performed the derivation of the $\bm{k}\cdot\bm{p}$ Hamiltonian using the Löwdin perturbation theory, and separately, using the folding-down approach. We also derived our effective Hamiltonian using the theory of invariants. The results obtained are in complete agreement with each other and differ from the model proposed in Ref. [70]. In addition, our investigation here goes beyond that of Ref. [70] by providing effective $\bm{k}\cdot\bm{p}$ parameters for the Na2CdSn DTI and deriving analytical expressions for both the energy dispersion and wave functions of the edge states arising in this material.
This paper is organized as follows. In Sec. II, we present the crystal structure and DFT calculations for Na2CdSn triatomic layers. Section III is devoted to the $\bm{k}\cdot\bm{p}$ model obtained via the Löwdin perturbation theory. We also present our results regarding the fitting of the effective model to the energy bands obtained via DFT. In Sec. IV, we analyze the topology of Na2CdSn, and in Sec. V, we calculate the analytical dispersion relations for the topological edge states. We also present our numerical study of the topological edge states in a ribbon geometry. We summarize our findings in Sec. VI.
II Na2CdSn triatomic layer: Crystal structure and DFT results
The Na2CdSn triatomic layer is constituted by Cd and Sn atoms forming a coplanar hexagonal lattice, while the sodium (Na) atoms lie above and below the center of each hexagon, forming a quasi-2D honeycomb material (see Fig. 1). This quasi-2D structure can be found as the constituting stacking block of the 3D Na2CdSn Dirac semimetal, already synthesized experimentally [46, 47]. In recent first-principles calculations, Mao et al. [70] have verified the stability of these compounds by performing phonon spectrum analysis and revealing only positive phonon energies.
[…]
@@ II Na2CdSn triatomic layer: Crystal structure and DFT results
In Fig. 1-a), we show a top view of the Na2CdSn. The coordinate system is chosen such that the $z$ axis points out of the plane of the figure. The primitive lattice vectors read $\bm{a}_{1}=\frac{a}{2}\left(\sqrt{3},1\right)$ and $\bm{a}_{2}=\left(0,a\right)$, where $a$ is the lattice parameter. In Fig. 1-b), we show a side view of the crystal structure with the observer looking from the positive $x$ axis. From this angle, the $x$ axis points out of the plane of the figure, and we can see more clearly the position of the Na atoms relative to Cd and Sn. The reciprocal lattice vectors can be shown to be $\bm{b}_{1}=\left(\frac{4\pi}{\sqrt{3}a},0\right)$ and $\bm{b}_{2}=\frac{2\pi}{\sqrt{3}a}\left(-1,\sqrt{3}\right)$. The first Brillouin zone (BZ) in this coordinate system is shown in Fig. 1-c), along with the high-symmetry points $\Gamma=\left(0,0\right)$, $M=\left(\frac{2\pi}{\sqrt{3}a},0\right)$ and $K=\frac{2\pi}{3a}\left(\sqrt{3},1\right)$ [71].
@@
The crystal structure of the 3D bulk compound belongs to the nonsymmorphic space group $D_{6h}^{4}$ ($P6_{3}/mmc$), with point group $D_{6h}$ ($6/mmm$) [72, 47, 45, 73, 74]. Our DFT results show that the Na2CdSn triatomic layer belongs to the $D_{3h}^{1}$ ($P\bar{6}m2$) space group, whose crystallographic point group is $D_{3h}$ ($\bar{6}m2$) [72, 47]. The corresponding set of generators are the $C_{3}$ rotation about the $z$ axis ($C_{3z}$), $C_{2}$ rotation about the $x$ axis ($C_{2x}$), and the mirror plane reflection about the $xy$ plane ($\sigma_{h}$) [see character table in the Supplemental Material (SM) [75] (see also Refs. [76, 77, 78] therein)].
@@
We investigate the electronic structure of Na2CdSn using DFT as implemented in the WIEN2k package [79], a full potential all-electron code employing the linearized augmented plane-wave plus local orbitals method. We use the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [80], a Monkhorst-Pack k-grid of 15$\times$15, and self-consistent convergence criteria of 10-6 $e$ (elementary charge) for the charge and 10-6 Ry for the energy. We consider a core–valence separation energy of $-6$ Ry, atomic spheres with orbital quantum numbers up to 10, and the plane-wave cutoff multiplied by the smallest atomic radii is set to 8. In the absence of spin-orbit coupling (SOC), we optimize the structural parameters and obtain an in-plane lattice parameter $a=4.978\,$Å and a thickness (vertical distance between Na atoms) of 3.3323.3323.332\,3.332Å (consistent with the lattice parameters of the 3D bulk compound [45, 46]). In the SM [75] we present the total energy obtained within DFT as a function of the lattice parameter $a$ and thickness $d$. In the presence of SOC, core electrons are considered fully relativistically, whereas valence electrons are treated in a second variational step [81], with the scalar-relativistic wave functions calculated in an energy window of $-10$ to $4$ Ry. We consider a vacuum region of 202020\,20Å.
@@
In Figs. 1-d) and 1-e), we show the electronic band structure and the orbital composition, respectively, obtained via DFT calculations for Na2CdSn in the absence of SOC. The Fermi level ($E_{F}$) lies at zero energy, and the blue dashed circle indicates the low-energy region we are interested in. The band structure is that of a semimetal with two $\Gamma_{6}$ bands fourfold degenerate at the $\Gamma$ point (accounting for the twofold spin degeneracy of each band). The closest remote bands belong to the $\Gamma_{1}$ and $\Gamma_{4}$ irreps [75]. We note that the $\Gamma_{6}$ bands are dominated by the $p_{x}$ and $p_{y}$ orbitals of the Cd and Sn atoms [dark blue in 1-e)], while the $\Gamma_{1}$ bands are dominated by $s$ orbitals (yellow) with a minor contribution from $p_{x}$ and $p_{y}$ (for more details, see SM [75]).
@@
In the presence of SOC, Fig. 1-f), the degeneracy at the $\Gamma$ point is lifted, and an energy gap $E_{g}=234.8\,$meV opens up. The material then becomes a semiconductor with topmost valence bands belonging to the $\Gamma_{7}$ and $\Gamma_{8}$ irreps, while the lowest conduction bands belong to $\Gamma_{7}$ and $\Gamma_{9}$. The next remote conduction band belongs to the $\Gamma_{8}$ irrep. All irreps are obtained from the total wave function.
@@
In Fig. 1-g), we show the expectation value of the $z$ component of spin (color coded) $\langle\hat{S}_{z}\rangle$ [$\langle\hat{S}_{z}\rangle>0$ ($\langle\hat{S}_{z}\rangle<0$) red (blue)]. We note that along the $\Gamma-M$ high-symmetry line ($\Sigma$), all energy bands are doubly degenerate. This high-symmetry line belongs to the point group $C_{2v}$, whose double group has only one irrep – the 2D $\Sigma_{5}$ [82]. The spin-up and down components of the electron wave functions transform according to the two basis functions of $\Sigma_{5}$. Therefore, all energy bands must be spin degenerate along this direction. Since any linear combination of the spin-degenerate states is also an eigenstate of the Hamiltonian, $\langle\hat{S}_{z}\rangle$ is not univocally defined along $\Gamma-M$. Hence, we represent the degenerate bands using the black color.
@@
The $\Gamma-K$ ($T$) and $M-K$ ($T^{\prime}$) lines, on the other hand, belong to the point group $C_{s}$, which consists of only two elements: the identity and the mirror symmetry $\sigma_{h}$. Its double group representation has two one-dimensional (1D) irreps, $T_{3}$ ($T^{\prime}_{3}$) and $T_{4}$ ($T^{\prime}_{4}$) for spin-up and -down states, respectively, along the $T$ ($T^{\prime}$) line. Thus, the energy bands are spin split along these directions. The spin expectation values of all eigenstates out of the $\Gamma-M$ line are given by $\langle\hat{S}_{z}\rangle\approx\pm\hbar/2$ and $\langle\hat{S}_{x}\rangle=\langle\hat{S}_{y}\rangle=0$, as required by mirror symmetry about the $z$ direction (see Appendix A) [54]. We note that our DFT calculated band structure is in great agreement with the previous report of Mao et al. [70].